OpenStudy (anonymous):

how do i find the midpoint when p1 to p2 is -1,4 ?

OpenStudy (anonymous):

Finding the midpoint is almost like finding the average. You add the absolute value of the two numbers. So 1+4. Now you divide that by 2. Which is 2.5. Now you add that to the smallest side. So your midpoint would be -1+2.5=1.5

OpenStudy (anonymous):

\[\frac{ x _{1} +x _{2}}{ 2} \] \[\frac{ y _{1} +y _{2}}{ 2} \] Just find the average and divide it by two. :)

OpenStudy (anonymous):



OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):